# Robustness of topological phases on aperiodic lattices

### v1

@@ 3.2.1. Position spectral triple over the Roe C*-algebra
is continuous for the norm topology on $\mathbb{B}(\mathcal{H}_{X})$. If $X$ does not contain discrete components, and $\mathcal{H}_{X}$ is the standard $X$-module as in Example 3.12, then the restriction of $\sigma$ to the $j$-th coordinate component of $\mathbb{R}^{d}$ gives the flow $\mathrm{e}^{\mathrm{i}t{\scriptstyle\mathsf{X}}_{j}}$ generated by ${\scriptstyle\mathsf{X}}_{j}$.
By conjugation, $\sigma\colon\mathbb{R}^{d}\to\mathbb{B}(\mathcal{H}_{X})$ induces a group homomorphism $\operatorname{Ad}\sigma\colon\mathbb{R}^{d}\to\operatorname{Aut}(\mathbb{B}(\mathcal{H}_{X}))$. We say that an operator $T\in\mathbb{B}(\mathcal{H}_{X})$ is continuous with respect to $\operatorname{Ad}\sigma$ iff the map $t\mapsto\operatorname{Ad}\sigma_{t}(T)$ is continuous for the norm topology on $\mathbb{B}(\mathcal{H})$. The property of being a norm limit of controlled operators can be described as a continuity property for the action $\operatorname{Ad}\sigma$:
@@
An operator $T\in\mathbb{B}(\mathcal{H}_{X})$ is a norm limit of controlled operators iff it is continuous with respect to $\operatorname{Ad}\sigma$. Thus $T\in\mathrm{C}^{*}_{\mathrm{Roe}}(X,\mathcal{H}_{X})$ iff it is locally compact and continuous with respect to $\operatorname{Ad}\sigma$.
We take a slight different viewpoint, interpreting the previous lemma as follows: the Roe C*-algebra is the largest C*-algebra, over which one can have a locally compact position spectral triple:
@@
$\left(\mathcal{A}\otimes\mathrm{C}\ell_{0,d},\quad\ell^{2}(\Lambda,\mathcal{K})\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d},\quad\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}\right)$
is a locally compact position spectral triple (cf. Definition 2.16), where $\mathcal{A}$ is a dense *-subalgebra of a C*-algebra $A$ represented on $\ell^{2}(\Lambda,\mathcal{K})$ by $\varphi\colon A\to\mathbb{B}(\ell^{2}(\Lambda,\mathcal{K}))$. Then $\varphi(\mathcal{A})$ is contained in the Roe C*-algebra $\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda,\mathcal{H}_{\Lambda})$ defined by the standard $\Lambda$-module $\mathcal{H}_{\Lambda}\mathrel{\vcentcolon=}\ell^{2}(\Lambda,\mathcal{K})$ as in Example 3.12.
@@
Identifying $A$ and $\mathcal{A}$ with their images under $\varphi$, we may assume that $\varphi\colon A\to\mathbb{B}(\ell^{2}(\Lambda,\mathcal{K}))$ is injective. Then we must show that every $a\in\mathcal{A}$ is locally compact and can be approximated by controlled operators. By definition of a locally compact spectral triple, it follows that $\left\langle x\,\middle|a\middle|\,\mathopen{}y\right\rangle\in\mathbb{K}(\mathcal{K})$ for all $x,y\in\Lambda$. This is the same condition for $a\in\mathcal{A}$ being locally compact in the sense of Definition 3.13.
Now we show that every $a\in\mathcal{A}$ is a norm limit of controlled operators. Since $[a\otimes\rho^{i},{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}]$ is bounded for all $a\in\mathcal{A}$ and $i,j\in\{1,\dots,d\}$, we have $[a,{\scriptstyle\mathsf{X}}_{j}]$ is a bounded operator for all $j\in\{1,\dots,d\}$, i.e. there exists $C>0$ such that $\left\lVert[a,{\scriptstyle\mathsf{X}}_{j}]\right\rVert\leq C$.
By the previous lemma, it suffices to show that for all $a\in\mathcal{A}$ and $j=1,\dots,d$, the map
$\mathbb{R}\to\mathbb{B}(\ell^{2}(\Lambda,\mathcal{K})),\qquad t\mapsto\mathrm{e}^{\mathrm{i}t{\scriptstyle\mathsf{X}}_{j}}a\mathrm{e}^{-\mathrm{i}t{\scriptstyle\mathsf{X}}_{j}}$
@@
$[\mathrm{e}^{it{\scriptstyle\mathsf{X}}_{j}},a]=\int_{0}^{1}\frac{\mathrm{d}}{\mathrm{d}s}\left(\mathrm{e}^{\mathrm{i}st{\scriptstyle\mathsf{X}}_{j}}a\mathrm{e}^{\mathrm{i}(1-s)t{\scriptstyle\mathsf{X}}_{j}}\right)\mathrm{d}s=\mathrm{i}t\int_{0}^{1}\mathrm{e}^{\mathrm{i}st{\scriptstyle\mathsf{X}}_{j}}[{\scriptstyle\mathsf{X}}_{j},a]\mathrm{e}^{\mathrm{i}(1-s)t{\scriptstyle\mathsf{X}}_{j}}\mathrm{d}s$
@@
$\left\lVert[\mathrm{e}^{\mathrm{i}t{\scriptstyle\mathsf{X}}_{j}},a]\right\rVert\leq t\int_{0}^{1}\left\lVert[{\scriptstyle\mathsf{X}}_{j},a]\right\rVert\mathrm{d}s\leq Ct.$
@@
$\left\lVert\mathrm{e}^{\mathrm{i}s{\scriptstyle\mathsf{X}}_{j}}a\mathrm{e}^{-\mathrm{i}s{\scriptstyle\mathsf{X}}_{j}}-\mathrm{e}^{\mathrm{i}t{\scriptstyle\mathsf{X}}_{j}}a\mathrm{e}^{-\mathrm{i}t{\scriptstyle\mathsf{X}}_{j}}\right\rVert$ $=\left\lVert\mathrm{e}^{\mathrm{i}t{\scriptstyle\mathsf{X}}_{j}}\left(\mathrm{e}^{\mathrm{i}(s-t){\scriptstyle\mathsf{X}}_{j}}a\mathrm{e}^{-\mathrm{i}(s-t){\scriptstyle\mathsf{X}}_{j}}-a\right)\mathrm{e}^{-\mathrm{i}t{\scriptstyle\mathsf{X}}_{j}}\right\rVert$
$=\left\lVert\mathrm{e}^{\mathrm{i}t{\scriptstyle\mathsf{X}}_{j}}[\mathrm{e}^{\mathrm{i}(s-t){\scriptstyle\mathsf{X}}_{j}},a]\mathrm{e}^{-\mathrm{i}(s-t){\scriptstyle\mathsf{X}}_{j}}\mathrm{e}^{-\mathrm{i}t{\scriptstyle\mathsf{X}}_{j}}\right\rVert$
$\leq\left\lVert[\mathrm{e}^{\mathrm{i}(s-t){\scriptstyle\mathsf{X}}_{j}},a]\right\rVert$
@@
Let $X\subseteq\mathbb{R}^{d}$ be a discrete countable set. Let $\mathcal{C}_{\mathrm{Roe}}(X)$ be the collection of operators $T\in\mathbb{B}(\ell^{2}(X,\mathcal{K}))$ satisfying:
@@
  • $\sup_{y\in X}\sum_{x\in X}\left\lVert T_{x,y}\right\rVert<+\infty$ and $\sup_{x\in X}\sum_{y\in X}\left\lVert T_{x,y}\right\rVert<+\infty$.
@@
(3.25) $\xi^{\mathrm{Roe}}_{X}\mathrel{\vcentcolon=}\left(\mathcal{C}_{\mathrm{Roe}}(X)\otimes\mathrm{C}\ell_{0,d},\quad\ell^{2}(X,\mathcal{K})\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d},\quad\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}\right)$
is a locally compact position spectral triple over $\mathrm{C}^{*}_{\mathrm{Roe}}(X)$. If $X$ is a Delone set, then $\xi^{\mathrm{Roe}}_{X}$ is $d$-summable.
@@
Let $T\in\mathcal{C}_{\mathrm{Roe}}(X)$, $x,y\in X$. Then
$\langle x\,|[T,{\scriptstyle\mathsf{X}}_{j}]|\,\mathopen{}y\rangle=\langle x\,|T{\scriptstyle\mathsf{X}}_{j}|\,\mathopen{}y\rangle-\langle x\,|{\scriptstyle\mathsf{X}}_{j}T|\,\mathopen{}y\rangle=(y_{j}-x_{j})T_{x,y}.$
@@
$\left\lVert[T,{\scriptstyle\mathsf{X}}_{j}]\right\rVert\leq\sup_{y\in X}\sum_{x\in X}R\cdot\left\lVert T_{x,y}\right\rVert<+\infty.$
This shows that $T$ maps the domain of ${\scriptstyle\mathsf{X}}_{j}$ into its domain and has bounded commutator with ${\scriptstyle\mathsf{X}}_{j}$ for all $j=1,\dots,d$.
Therefore, for any $S\in\mathrm{C}\ell_{0,d}$, $T\otimes S$ maps the domain of $\sum_{j=1}^{n}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}$ into the domain of $\sum_{j=1}^{n}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}$ and
$\left[T\otimes S,\quad\sum_{j=1}^{n}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}\right]=\sum_{j=1}^{n}\left[T,{\scriptstyle\mathsf{X}}_{j}\right]\otimes\left[S,\gamma^{j}\right]$
is bounded. The operator
$\left(1+\left(\sum_{j=1}^{n}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}\right)^{2}\right)^{-1}=\left(1+\sum_{j=1}^{n}{\scriptstyle\mathsf{X}}_{j}^{2}\otimes 1\right)^{-1}$
is compact. Therefore, $\xi^{\mathrm{Roe}}_{X}$ is a spectral triple, which is a locally compact position spectral triple because $T$ is locally compact.
If $X$ is a Delone set, then it is uniformly discrete, which implies that
$\operatorname{tr}\left(1+\left(\sum_{j=1}^{n}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}\right)^{2}\right)^{-s/2}=2^{d}\sum_{x\in X}\left(1+\lvert x\rvert^{2}\right)^{-s/2}$
is finite for all $s>d$. ∎
3.2.2. Position spectral triple generates K-homology
Towards the end of this section, we shall prove that the position spectral triple $\xi_{\Lambda}^{\mathrm{Roe}}$ induces an isomorphism
$\operatorname{KK}(\mathrm{C}\ell_{d,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda))\xrightarrow{\sim}\operatorname{KK}(\mathrm{C}\ell_{d,d},\mathbb{R})\simeq\mathbb{Z}.$
The proof is based on these observations.
@@ 3.2.2. Position spectral triple generates K-homology
    If $\Lambda=\mathbb{Z}^{d}$, then the K-homology class of the position spectral triple $\xi_{\mathbb{Z}^{d}}^{\mathrm{Roe}}$ pulls back to Kasparov’s Dirac element for the “real” manifold $\mathbb{T}^{d}$ which induces an isomorphism $\operatorname{KK}(\mathrm{C}\ell_{d,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d}))\xrightarrow{\sim}\mathbb{Z}$;
@@
    Any Delone set $\Lambda\subseteq\mathbb{R}^{d}$ is coarsely equivalent to $\mathbb{Z}^{d}$, hence $\xi^{\mathrm{Roe}}_{\Lambda}$ induces the same map as $\xi^{\mathrm{Roe}}_{\mathbb{Z}^{d}}$ up to an isomorphism of the K-theory groups
    $\operatorname{KK}(\mathrm{C}\ell_{d,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d}))\simeq\operatorname{KK}(\mathrm{C}\ell_{d,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)).$
@@
(3.27) $\xi^{\mathrm{Roe}}_{\mathbb{Z}^{d}}\mathrel{\vcentcolon=}\left(\mathcal{C}_{\mathrm{Roe}}(\mathbb{Z}^{d})\otimes\mathrm{C}\ell_{0,d},\quad\ell^{2}(\mathbb{Z}^{d},\mathcal{K})\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d},\quad\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\gamma^{j}\right),$
induces an isomorphism
$\operatorname{KK}(\mathrm{C}\ell_{d,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d}))\xrightarrow{\sim}\operatorname{KK}(\mathrm{C}\ell_{d,d},\mathbb{R})\simeq\mathbb{Z}.$
The proof below comes from [Ewert-Meyer:Coarse_geometry]*Section 4, which we recall for the reader’s convenience. Here we must use “real” C*-algebras and KKR-theory to describe the Dirac element of the “real” manifold $\mathbb{T}^{d}$. While passing to the “real” group C*-algebras of $\mathbb{Z}^{d}$ by Fourier transform, these $\operatorname{KKR}$-theory groups simply to $\operatorname{KKO}$-theory.

### v2

@@ 3.2.1. Position spectral triple over the Roe C*-algebra
is continuous for the norm topology on $\mathbb{B}(\mathcal{H}_{X})$. If $X$ does not contain discrete components, and $\mathcal{H}_{X}$ is the standard $X$-module as in Example˜3.15, then the restriction of $\sigma$ to the $j$-th coordinate component of $\mathbb{R}^{d}$ gives the flow $\mathrm{e}^{\mathrm{i}t{\scriptstyle\mathsf{X}}_{j}}$ generated by ${\scriptstyle\mathsf{X}}_{j}$.
By conjugation, $\sigma\colon\mathbb{R}^{d}\to\mathbb{B}(\mathcal{H}_{X})$ induces a group homomorphism $\operatorname{Ad}\sigma\colon\mathbb{R}^{d}\to\operatorname{Aut}(\mathbb{B}(\mathcal{H}_{X}))$. We say that an operator $T\in\mathbb{B}(\mathcal{H}_{X})$ is continuous with respect to $\operatorname{Ad}\sigma$ if and only if the map $t\mapsto\operatorname{Ad}\sigma_{t}(T)$ is continuous for the norm topology on $\mathbb{B}(\mathcal{H})$. The property of being a norm limit of controlled operators can be described as a continuity property for the action $\operatorname{Ad}\sigma$:
@@
An operator $T\in\mathbb{B}(\mathcal{H}_{X})$ is a norm limit of controlled operators if and only if it is continuous with respect to $\operatorname{Ad}\sigma$. Thus $T\in\mathrm{C}^{*}_{\mathrm{Roe}}(X,\mathcal{H}_{X})$ if and only if it is locally compact and continuous with respect to $\operatorname{Ad}\sigma$.
We take a slight different viewpoint, interpreting the previous lemma as follows: the Roe C*-algebra is the largest C*-algebra, over which one can have a position spectral triple:
@@
$\xi\mathrel{\vcentcolon=}\left(\mathcal{A}\otimes\mathrm{C}\ell_{0,d},\quad\ell^{2}(\Lambda)\otimes\mathcal{K}\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d},\quad\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\operatorname{id}_{\mathcal{K}}\otimes\gamma^{j}\right)$
is a position spectral triple (cf. Definition˜2.11), where $\mathcal{A}$ is a dense *-subalgebra of a C*-algebra $A$ represented on $\ell^{2}(\Lambda)\otimes\mathcal{K}$ by $\varphi\colon A\to\mathbb{B}(\ell^{2}(\Lambda)\otimes\mathcal{K})$. Then $\varphi(\mathcal{A})$ is contained in the Roe C*-algebra $\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda,\mathcal{H}_{\Lambda})$ defined by the standard $\Lambda$-module $\mathcal{H}_{\Lambda}\mathrel{\vcentcolon=}\ell^{2}(\Lambda)\otimes\mathcal{K}$ as in Example˜3.15.
@@
Identifying $A$ and its dense subalgebra $\mathcal{A}$ with their images under $\varphi$, we may assume that $\varphi\colon A\to\mathbb{B}(\ell^{2}(\Lambda)\otimes\mathcal{K})$ is injective. Then we must show that every $a\in\mathcal{A}$ is locally compact and can be approximated by controlled operators. Being locally compact amounts to saying that $\langle x\,|\varphi(a)|\,\mathopen{}y\rangle\in\mathbb{K}(\mathcal{K})$ for all $x,y\in\Lambda$, which always holds by Lemma˜2.13 (providing that $\xi$ is a position spectral triple.)
Now we show that every $a\in\mathcal{A}$ is a norm limit of controlled operators. For convenience, below we write $\hat{{\scriptstyle\mathsf{X}}}_{j}$ for the operator ${\scriptstyle\mathsf{X}}_{j}\otimes\operatorname{id}_{\mathcal{K}}$ on $\ell^{2}(\Lambda)\otimes\mathcal{K}$.
Since $a\otimes\rho^{i},\hat{{\scriptstyle\mathsf{X}}}_{j}\otimes\gamma^{j}]$ is bounded for all $a\in\mathcal{A}$ and $i,j\in\{1,\dots,d\}$, we have $a,\hat{{\scriptstyle\mathsf{X}}}_{j}]$ is a bounded operator for all $j\in\{1,\dots,d\}$, i.e. there exists $C>0$ such that $\left\lVert[a,\hat{{\scriptstyle\mathsf{X}}}_{j}]\right\rVert\leq C$.
By Lemma˜3.25, it suffices to show that for all $a\in\mathcal{A}$ and $j=1,\dots,d$, the map
$\mathbb{R}\to\mathbb{B}(\ell^{2}(\Lambda,\mathcal{K})),\qquad t\mapsto\mathrm{e}^{\mathrm{i}t\hat{{\scriptstyle\mathsf{X}}}_{j}}a\mathrm{e}^{-\mathrm{i}t\hat{{\scriptstyle\mathsf{X}}}_{j}}$
@@
$\mathrm{e}^{it\hat{{\scriptstyle\mathsf{X}}}_{j}},a]=\int_{0}^{1}\frac{\mathrm{d}}{\mathrm{d}s}\left(\mathrm{e}^{\mathrm{i}st\hat{{\scriptstyle\mathsf{X}}}_{j}}a\mathrm{e}^{\mathrm{i}(1-s)t\hat{{\scriptstyle\mathsf{X}}}_{j}}\right)\mathrm{d}s=\mathrm{i}t\int_{0}^{1}\mathrm{e}^{\mathrm{i}st\hat{{\scriptstyle\mathsf{X}}}_{j}}[\hat{{\scriptstyle\mathsf{X}}}_{j},a]\mathrm{e}^{\mathrm{i}(1-s)t\hat{{\scriptstyle\mathsf{X}}}_{j}}\mathrm{d}s$
@@
$\left\lVert[\mathrm{e}^{\mathrm{i}t\hat{{\scriptstyle\mathsf{X}}}_{j}},a]\right\rVert\leq t\int_{0}^{1}\left\lVert[\hat{{\scriptstyle\mathsf{X}}}_{j},a]\right\rVert\mathrm{d}s\leq Ct.$
@@
$\left\lVert\mathrm{e}^{\mathrm{i}s\hat{{\scriptstyle\mathsf{X}}}_{j}}a\mathrm{e}^{-\mathrm{i}s\hat{{\scriptstyle\mathsf{X}}}_{j}}-\mathrm{e}^{\mathrm{i}t\hat{{\scriptstyle\mathsf{X}}}_{j}}a\mathrm{e}^{-\mathrm{i}t\hat{{\scriptstyle\mathsf{X}}}_{j}}\right\rVert$ $=\left\lVert\mathrm{e}^{\mathrm{i}t\hat{{\scriptstyle\mathsf{X}}}_{j}}\left(\mathrm{e}^{\mathrm{i}(s-t)\hat{{\scriptstyle\mathsf{X}}}_{j}}a\mathrm{e}^{-\mathrm{i}(s-t)\hat{{\scriptstyle\mathsf{X}}}_{j}}-a\right)\mathrm{e}^{-\mathrm{i}t\hat{{\scriptstyle\mathsf{X}}}_{j}}\right\rVert$
$=\left\lVert\mathrm{e}^{\mathrm{i}t\hat{{\scriptstyle\mathsf{X}}}_{j}}[\mathrm{e}^{\mathrm{i}(s-t)\hat{{\scriptstyle\mathsf{X}}}_{j}},a]\mathrm{e}^{-\mathrm{i}(s-t)\hat{{\scriptstyle\mathsf{X}}}_{j}}\mathrm{e}^{-\mathrm{i}t\hat{{\scriptstyle\mathsf{X}}}_{j}}\right\rVert$
$\leq\left\lVert[\mathrm{e}^{\mathrm{i}(s-t)\hat{{\scriptstyle\mathsf{X}}}_{j}},a]\right\rVert$
@@
Let $\Lambda\subseteq\mathbb{R}^{d}$ be a discrete countable set. Let $\mathcal{C}_{\mathrm{Roe}}(\Lambda)$ be the collection of operators $T\in\mathbb{B}(\ell^{2}(\Lambda,\mathcal{K}))$ satisfying:
@@
  • $\sup_{y\in\Lambda}\sum_{x\in\Lambda}\left\lVert T_{x,y}\right\rVert<+\infty$ and $\sup_{x\in\Lambda}\sum_{y\in\Lambda}\left\lVert T_{x,y}\right\rVert<+\infty$.
@@
(3.28) $\xi^{\mathrm{Roe}}_{\Lambda}\mathrel{\vcentcolon=}\left(\mathcal{C}_{\mathrm{Roe}}(\Lambda)\otimes\mathrm{C}\ell_{0,d},\quad\ell^{2}(\Lambda)\otimes\mathcal{K}\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d},\quad\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\operatorname{id}_{\mathcal{K}}\otimes\gamma^{j}\right)$
is a position spectral triple over $\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)$.
@@
Let $T\in\mathcal{C}_{\mathrm{Roe}}(\Lambda)$, $x,y\in\Lambda$. Then
$\langle x\,|[T,{\scriptstyle\mathsf{X}}_{j}\otimes\operatorname{id}_{\mathcal{K}}]|\,\mathopen{}y\rangle=\langle x\,|T({\scriptstyle\mathsf{X}}_{j}\otimes\operatorname{id}_{\mathcal{K}})|\,\mathopen{}y\rangle-\langle x\,|({\scriptstyle\mathsf{X}}_{j}\otimes\operatorname{id}_{\mathcal{K}})T|\,\mathopen{}y\rangle=(y_{j}-x_{j})T_{x,y}.$
@@
$\left\lVert[T,{\scriptstyle\mathsf{X}}_{j}\otimes\operatorname{id}_{\mathcal{K}}]\right\rVert\leq\sup_{y\in\Lambda}\sum_{x\in\Lambda}R\cdot\left\lVert T_{x,y}\right\rVert<+\infty.$
So $T$ preserves the domain of ${\scriptstyle\mathsf{X}}_{j}\otimes\operatorname{id}_{\mathcal{K}}$ and boundedly commutes with it for every $j=1,\dots,d$.
Therefore, for any $S\in\mathrm{C}\ell_{0,d}$, $T\otimes S$ preserves the domain of ${\scriptstyle\mathsf{X}}\mathrel{\vcentcolon=}\sum_{j=1}^{n}{\scriptstyle\mathsf{X}}_{j}\otimes\operatorname{id}_{\mathcal{K}}\otimes\gamma^{j}$, and boundedly commutes with $\scriptstyle\mathsf{X}$}. The local compactness of $T$ implies that the operator
$T\left(1+\left(\sum_{j=1}^{n}{\scriptstyle\mathsf{X}}_{j}\otimes\operatorname{id}_{\mathcal{K}}\otimes\gamma^{j}\right)^{2}\right)^{-1}=T\left(1+\sum_{j=1}^{n}{\scriptstyle\mathsf{X}}_{j}^{2}\right)^{-1}\otimes\operatorname{id}_{\mathcal{K}}\otimes\operatorname{id}_{\bigwedge\nolimits^{*}\mathbb{R}^{d}}$
is compact. Therefore, $\xi^{\mathrm{Roe}}_{\Lambda}$ is a position spectral triple. ∎
3.2.2. Position spectral triple generates $\operatorname{KO}_{*}(\mathbb{R})$
Towards the end of this section, we shall prove that the position spectral triple $\xi_{\Lambda}^{\mathrm{Roe}}$ induces isomorphisms
$\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda))\xrightarrow{\sim}\operatorname{KK}(\mathrm{C}\ell_{n,d},\mathbb{R})$
for all $n$. The proof is based on these observations.
@@ 3.2.2. Position spectral triple generates $\operatorname{KO}_{*}(\mathbb{R})$
    If $\Lambda=\mathbb{Z}^{d}$, then the KK-class of the position spectral triple
    $\xi^{\mathrm{Roe}}_{\mathbb{Z}^{d}}]\in\operatorname{KK}(\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d})_{\mathbb{R}}\otimes\mathrm{C}\ell_{0,d},\mathbb{R})$
    pulls back to the fundamental class of $\operatorname{KK}(\mathrm{C}^{*}(\mathbb{Z}^{d})_{\mathbb{R}}\otimes\mathrm{C}\ell_{0,d},\mathbb{R})$ as in Theorem˜2.15. Therefore, for each $n$, it induces an isomorphism
    $\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d}))\xrightarrow{\sim}\operatorname{KK}(\mathrm{C}\ell_{n,d},\mathbb{R}).$
@@
    Any other Delone set $\Lambda\subseteq\mathbb{R}^{d}$ is coarsely equivalent to $\mathbb{Z}^{d}$. Therefore, $\xi^{\mathrm{Roe}}_{\Lambda}$ induces the same map as $\xi^{\mathrm{Roe}}_{\mathbb{Z}^{d}}$ up to the isomorphism given by
    $\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d}))\simeq\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\Lambda)).$
@@
(3.30) $\xi^{\mathrm{Roe}}_{\mathbb{Z}^{d}}\mathrel{\vcentcolon=}\left(\mathcal{C}_{\mathrm{Roe}}(\mathbb{Z}^{d})\otimes\mathrm{C}\ell_{0,d},\quad\ell^{2}(\mathbb{Z}^{d})\otimes\mathcal{K}\otimes\bigwedge\nolimits^{*}\mathbb{R}^{d},\quad\sum_{j=1}^{d}{\scriptstyle\mathsf{X}}_{j}\otimes\operatorname{id}_{\mathcal{K}}\otimes\gamma^{j}\right),$
induces, for each $n$, an isomorphism
$\operatorname{KK}(\mathrm{C}\ell_{n,0},\mathrm{C}^{*}_{\mathrm{Roe}}(\mathbb{Z}^{d}))\xrightarrow{\sim}\operatorname{KK}(\mathrm{C}\ell_{n,d},\mathbb{R}).$